OpenStudy (anonymous):

Find the ratio of the sum & product of the roots 7x^2 - 12x + 18 = 0.

Parth (parthkohli):

Use Vieta's Formula

OpenStudy (anonymous):

sum of roots = -b/a

Parth (parthkohli):

Yea

Parth (parthkohli):

Product is \(\dfrac{c}{a}\)

OpenStudy (anonymous):

vat's ur answer cmg is it 2/3

Parth (parthkohli):

You could just solve the equation then.

OpenStudy (anonymous):

Yup..u r correct

OpenStudy (anonymous):

@msingh

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

thanks @Yahoo!

OpenStudy (anonymous):

Welcome..;)

Parth (parthkohli):

\[\dfrac{\dfrac{-b}{a}}{\dfrac{c}{a}} = \dfrac{-b}{c}\]So that's why.